# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Appendix B Threading flux
When dimensionally-reducing the lattice system to the cylinder, the charge and spin flux through the cross sections of the cylinder distinguish gauge-inequivalent systems. This additional flux can be implemented by adding a constant term to the vector potential:
@@
For the purposes of numerical implementation, particularly in hybrid space where we exchange $y$ for $k_{y}$, it is convenient to instead implement flux threading via a twist in the boundary conditions. Starting with periodic boundary conditions $c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r})$ for $\phi=0$, define a new set of operators
@@
Then instead of implementing Eq. (24), we can instead replace every instance of $c$ with $c_{\phi}$. This does not affect terms that depend solely on charge and spin densities, e.g., the Hubbard interaction. However, it will modify the hoppings in a way that is equivalent to Eq. (24). Under the operator replacement, the hoppings are replaced by
@@
The utility of this alternative formulation is that it leads to a clean numerical implementation upon fourier transforming $y\to k_{y}$. In particular, recall that $(x,y)=(\bm{r}\cdot\bm{b}_{1}/2\pi,\bm{r}\cdot\bm{b}_{2}/2\pi)$ and define
@@
Note that this is the Fourier convention of TeNPy. When we express the Hamiltonian in the hybrid $xk_{y}$ representation, we find that all that’s required is to shift the allowed values of $k_{y}.$ To see this, we impose the twisted boundary condition Eq. (33) on the Fourier-expanded operators:
@@
The allowed values of momenta are therefore quantized to:
@@
which are shifted relative to the $\phi=0$ case. When the original Hamiltonian $\hat{H}_{\phi=0}$ is written in hybrid space, some coefficients will depend on $k.$ The only affect of the replacement Eq. (29) is then for those coefficients to instead be evaluated on these shifted set of allowed values.
@@ Appendix C Two-band hopping Hamiltonian
Now let us specialize to the case $(p,q)=(1,2)$, which corresponds to $\Phi_{\triangle}=\pi/2$. To specify the unit cells (which each contain two sites), we often write $x=\tau+2u$, where $u\in\mathbb{Z}$, while $\tau\in\{0,1\}$ specifies the even/odd rings. With periodic boundary conditions and before flux insertion, the hoppings Eq. (15) take the particular form
@@
$=t\exp\left(-i\pi(x+1/2)\right).$ (40)
We depict the resulting hopping phases in Fig. S1(a) for the lattice orientation where $\bm{a}_{2}=a\hat{y}$ is a nearest-neighbour bond; this becomes becomes the YC geometry when the system is placed on a cylinder under the identification $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$.
Figure S1: (a) Hopping phases for the $(p,q)=(1,2)$ two-band Hamiltonian corresponding to $\pi/2$ flux per triangle. The two black circles indicate the two sites in the magnetic unit cell. The phases indicated by the color bar are accrued when hopping upward, right-upward, and right-downward; the complex-conjugate phases are accrued when hopping against these directions. (b) Graphical depiction of the fluxes through various plaquettes upon placing the lattice on a cylinder. Every triangle on the surface of the infinite cylinder is penetrated by $\Phi_{\triangle}=\pi/2$ external flux. As a consequence of Gauss’s law, the external flux through adjacent rings—each of which is a polygonal surface bounded by a ring of length $L_{y}$—must alternate by $\pi$ when $L_{y}$ is odd, and must be uniform when $L_{y}$ is even.
@@
This affects the hopping directions with some weight along $\bm{\delta}=\bm{a}_{2}$ and, by comparison to Eq. (28) above, corresponds to threading $\phi=\pi L_{y}$ through the cylinder. In the 2D plane geometry — and later on some even-$L_{y}$ geometries as we’ll show — the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation. In particular, define the magnetic translation operation by
$T_{x}^{\text{mag}}c_{\bm{r}}^{\dagger}(T_{x}^{\text{mag}})^{-1}=c_{\bm{r}+\bm{a}_{1}}^{\dagger}\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)=c_{\bm{r}+\bm{a}_{1}}^{\dagger}(-1)^{y},$ (48)
@@
$T_{x}^{\text{mag}}\hat{h}(T_{x}^{\text{mag}})^{-1}$ $=T_{x}^{\text{mag}}\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(T_{x}^{\text{mag}})^{-1}$ (49)
@@
Consider the naive definition of the particle-hole operation:
@@
where the condition specifies it is a linear (as opposed to anti-linear) unitary many-body operator. If 1 and 2 are an adjacent pair of sites, then particle-hole acts as
@@
so that it flips the Peierl’s phase and increases it by $\pi.$ On the triangular lattice, this modifies the flux through each triangle as follows:
@@
Requiring this operation to be invariant mod $2\pi,$ we find that we need
$\Phi_{\text{par}}=2\Phi_{\triangle}=\pi\left(2n+1\right),\qquad n\in\mathbb{Z}.$ (59)
This is consistent with the observation that the Hofstadter spectrum on the triangular lattice only has bands with opposite energy for these values of $\Phi_{\triangle}$. In particular, in this work we consider $\Phi_{\triangle}=\pi/2$. We find at each momentum $\bm{k}$ that both the energy and Berry curvature are exactly opposite, with the bands having Chern number ±$\pm 1.$
In general, the naive particle-hole must be composed with a gauge transformation in order to restore all the hopping elements. Given the particular form of the hoppings Eq. (SI) to Eq. (SI) for the two-band model, we obtain the following form of the particle-hole symmetry:
@@
where $\mathsf{T}$ is the transposition operation on the row vector of electronic annihilation operators.
SII.3 Mirror composed with time-reversal
The orbital magnetic field breaks the time-reversal and mirror symmetries enjoyed by the $\Phi_{\triangle}\equiv 0$ (mod $\pi$) Hubbard models on the triangular lattice. However, the combination of mirror with time-reversal remains a symmetry and should not be spontaneously broken by the CSL or IQH phases. In particular, we focus on $M_{y}\mathcal{T}$ since this symmetry acts “on-site” (i.e., within each ring) when the system is placed on the cylinder. Recall that $M_{y}$ should send $\bm{r}\to M_{y}\bm{r}$ (flipping the Cartesian $y$-coordinate) while also mapping $\bm{S}\to(-S_{x},S_{y},-S_{z}).$ On the other hand, time-reversal should flip all components of spin. Their combination should therefore map $\bm{S}\to(S_{x},-S_{y},S_{z})$, which is easy to implement due to $\mathcal{T}$ being anti-unitary. We claim that the desired symmetry is
$(M_{y}\mathcal{T})c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-1}=c^{\dagger}(M_{y}\bm{r})e^{i(-1)^{x}\pi/4},$ (61)
where the exponential factor is chosen to restore the phases of the hoppings. We confirm this by explicit calculation:
$(M_{y}\mathcal{T})\hat{h}(M_{y}\mathcal{T})^{-1}$ $=(M_{y}\mathcal{T})\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(M_{y}\mathcal{T})$ (62)
$=\sum c_{M_{y}(\bm{r}+\bm{\delta})}^{\dagger}e^{i(-1)^{\delta_{x}+x}\pi/4}t^{*}_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}e^{-i(-1)^{x}\pi/4}c_{M_{y}\bm{r}}$ (63)
$=\sum c_{\bm{r}+\bm{\delta}}^{\dagger}e^{i(-1)^{x}\left((-1)^{\delta_{x}}-1\right)\pi/4}t^{*}_{M_{y}(\bm{r}+\bm{\delta})\leftarrow M_{y}\bm{r}}c_{\bm{r}},$ (64)
which can be shown to be respected for Eqs. (SI) to (SI) above.
It is easy to verify that $M_{y}\mathcal{T}$ is not Kramers. However, it can be convenient to define a Kramers symmetry by performing an additional $\pi$ spin rotation of about the $y$ axis:
$(M_{y}\mathcal{T}^{\prime})c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}=c^{\dagger}(M_{y}\bm{r})(i\sigma_{y})e^{i(-1)^{x}\pi/4}.$ (65)
One then has that
$(M_{y}\mathcal{T})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-2}$ $=+c^{\dagger}(\bm{r}),$ (66)
$(M_{y}\mathcal{T}^{\prime})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-2}$ $=-c^{\dagger}(\bm{r}).$ (67)
Furthermore, we note that
$(M_{y}\mathcal{T})\bm{S}(\bm{r})(M_{y}\mathcal{T})^{-1}$ $=\left(S_{x},-S_{y},S_{z}\right)_{M_{y}\bm{r}},$ (68)
$(M_{y}\mathcal{T}^{\prime})\bm{S}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}$ $=-\bm{S}(M_{y}\bm{r}).$ (69)
SIII Symmetries on the cylinder
To implement charge-flux threading, we work in the real-space formalism with periodic boundary conditions and vector potential modified according to Eq. (24). We then get
$t_{(x,y+1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(i\pi x\right)\exp\left(i\phi/L_{y}\right)$ (70)
$t_{(x+1,y)\leftarrow(x,y)}^{\phi}$ $=t$ (71)
$t_{(x+1,y-1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(-i\pi(x+1/2)\right)\exp\left(-i\phi/L_{y}\right),$ (72)
so that we pick up an additional phase of $\phi/L_{y}$ when $y$ increments by $1$ One finds that the symmetries of the system are sensitive not only to the parity of $L_{y}$, but also to the threaded charge and spin flux.
SIII.1 Magnetic translations
Figure S2: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively. The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
When $L_{y}$ is even, the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation. In contrast, by Fig. S1(b), we know that the effect of translations on the fluxes through loops cannot be gauged away when $L_{y}$ is odd. This manifests in the fact that $T_{x}^{\text{mag}}$ [see Eq. (48)] is ill-defined, which we see by imposing periodic boundary conditions and comparing $\bm{r}+L_{y}\bm{a}_{2}$ with $\bm{r}$; note that
$\exp\left(\frac{i}{2}\bm{b}_{2}\cdot(\bm{r}+L_{y}\bm{a}_{2})\right)=\exp\left(i\pi L_{y}\right)\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)\neq\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right).$ (73)
SIII.2 Particle-hole
Now, let’s fix $\Phi_{\text{par}}=2\Phi_{\triangle}=+\pi$ and interrogate whether placing the system on a YC cylinder changes this conclusion. Focusing on a particular ring $r$ around the cylinder with flux $\Phi_{r},$ we note that
@@
Therefore, $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to a gauge transformation) when
@@
for all rings $r.$
When $L_{y}$ is even, Gauss’s law dictates that the flux $\Phi_{r_{1,2}}$ through two adjancent rings $r_{1,2}$ must be equal — note that this allows $T_{x}$ to be a symmetry (up to a gauge transformation). Therefore, possibly after threading $\phi$ external flux through each ring of the cylinder, it is possible to tune all the $\Phi_{r}$ simultaneously (to 00 or $\pi$) so that $\tilde{\mathcal{\mathcal{P}}}$ is respected. Only in those cases is $T_{x}\tilde{\mathcal{\mathcal{P}}}$ also a symmetry (up to gauge).
@@
This always rules out $T_{x}$ as a symmetry. However, so long as $\Phi_{r_{1}}$ is $\pi/2$ or $-\pi/2,$ both Eqs. (75) and (76) can be respected, so that $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge). However, odd $L_{y}$ is special in that if we tune $\Phi_{r_{1}}$ to 0 or $\pi,$ then $T_{x}\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge) even though neither $T_{x}$ nor $\tilde{\mathcal{\mathcal{P}}}$ are themselves symmetries.
Appendix D Observable and Correlation Functions
SI Heisenberg Operator in terms of Fermions
If the sites are spinless FermionSites and up and down spins are encoded in adjacent FermionSites, we can compute spin correlations indirectly using the fermion creation and annihilation operators. Since the spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ involves the product of two spin operators, we need to express it in terms of the fermion operators. In a system with spinless FermionSites, we can define the spin operators as:
$S_{i}^{+}=c_{i\uparrow}^{\dagger}c_{i\downarrow},\quad S_{i}^{-}=c_{i\downarrow}^{\dagger}c_{i\uparrow},\quad S_{i}^{z}=\frac{1}{2}(c_{i\uparrow}^{\dagger}c_{i\uparrow}-c_{i\downarrow}^{\dagger}c_{i\downarrow}),$ (77)

### v2

@@ Appendix B Threading flux
When dimensionally-reducing the lattice system to the cylinder, the charge and spin flux through the cross sections of the cylinder distinguish gauge-inequivalent systems. This additional flux can be implemented by adding a constant offset to the vector potential:
@@
For the purposes of numerical implementation, particularly in hybrid space where we trade $y$ for $k_{y}$, it is convenient to instead implement flux threading via a twist in the boundary conditions. Starting with periodic boundary conditions for the fermion operators at $\phi=0$, namely $c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r})$, we define a new set of operators
@@
Then instead of implementing Eq. (25), we can instead replace every instance of $c$ in the second-quantized interacting Hamiltonian with $c_{\phi}$. This does not affect terms that depend solely on charge and spin densities, e.g., the Hubbard interaction. However, it modifies the hoppings in a way that is equivalent to Eq. (25). Under the operator replacement, the hoppings are replaced by
@@
The utility of this alternative formulation is that it leads to a clean numerical implementation upon fourier transforming $y\to k_{y}$. In particular, note that $(x,y)=(\bm{r}\cdot\bm{b}_{1}/2\pi,\bm{r}\cdot\bm{b}_{2}/2\pi)$ and define
@@
Note that this is the Fourier convention of TeNPy. When we express the Hamiltonian in the hybrid $xk_{y}$ representation, we find that all that is required is to shift the allowed values of $k_{y}.$ To see this, we impose the twisted boundary condition Eq. (34) on the Fourier-expanded operators:
@@
The allowed values of momenta are therefore quantized to
@@
which are shifted relative to the $\phi=0$ case. When the original Hamiltonian $\hat{H}_{\phi=0}$ is written in hybrid space, some coefficients will depend on $k.$ The only effect of the replacement Eq. (30) is then for those coefficients to instead be evaluated on this shifted set of allowed values [8].
@@ Appendix C Two-band hopping Hamiltonian
Figure S2: (a) Plot of the lower and upper energy bands of the Hofstadter system with $(p,q)=(1,2)$, corresponding to the case with $\Phi_{\triangle}=\pi/2$ considered throughout the main text. Momenta are given within the first “magnetic” Brillouin zone, defined with respect to the magnetic reciprocal vectors of Eq. (7) in the 2D planar YC coordinate system of Eq. (2). (b) Plot of the Berry curvature of the aforementioned bands which, like their energy, is equal and opposite at every momentum. The bands carry opposite Chern numbers $\pm 1$.
@@
Now let us specialize to the YC lattice geometry and to the case $(p,q)=(1,2)$, which corresponds to $\Phi_{\triangle}=\pi/2$. To specify the unit cells (which each contain two sites), we write $x=\tau+2u$, where $u\in\mathbb{Z}$, while $\tau\in\{0,1\}$ specifies the even/odd rings. With periodic boundary conditions and absent any flux insertion, the hoppings Eq. (16) take the particular form
@@
$=t\exp\left(-i\pi(x+1/2)\right),$ (41)
as shown graphically in Fig. S1(a). We depict the resulting energy bands $E_{\text{lower/upper}}(\bm{k})$ in Fig. S2, as well as their corresponding Berry curvature $\Omega_{\text{lower/upper}}(\bm{k})$, in the YC coordinate system consistent with Eqs. (2) and (7). The energies are equal and opposite at every momentum and are gapped everywhere in the Brillouin zone. The Berry curvature is also everywhere equal and opposite, with the Chern numbers evaluating to +1+1 for the lower band and 1-1 for the upper band.
@@
This affects the hoppings in directions with some weight along $\bm{\delta}=\bm{a}_{2}$ and, by comparison to Eq. (29) above, corresponds to threading $\phi=\pi L_{y}$ through the cylinder. In the 2D plane geometry—and on cylinder geometries of even circumference as we’ll later show—the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation. In particular, define the magnetic translation operation by
$T_{x}c_{\bm{r}}^{\dagger}T_{x}^{-1}=c_{\bm{r}+\bm{a}_{1}}^{\dagger}\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)=c_{\bm{r}+\bm{a}_{1}}^{\dagger}(-1)^{y},$ (49)
@@
$T_{x}\hat{h}T_{x}^{-1}$ $=T_{x}\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)T_{x}^{-1}$ (50)
@@
Consider the most naive definition of the particle-hole operation:
@@
where the second condition specifies that it is a linear (as opposed to anti-linear) unitary many-body operator. If 1 and 2 denote an adjacent pair of sites, then this naive particle-hole operation acts as
@@
so that it reverses the Peierl’s phase and shifts it by $\pi.$ On the triangular lattice, this modifies the flux through each elementary triangular plaquette as follows:
@@
Requiring that this operation leave the flux invariant modulo $2\pi,$ we find that we need
$2\Phi_{\triangle}=\pi\left(2n+1\right),\qquad n\in\mathbb{Z}.$ (60)
This is consistent with the observation that the Hofstadter spectrum on the triangular lattice only has bands with opposite energy for these values of $\Phi_{\triangle}$. In particular, in this work we consider $\Phi_{\triangle}=\pi/2$. We find at each momentum $\bm{k}$ that both the energy and Berry curvature are exactly opposite, with the bands having Chern number $\pm 1.$
For a generic choice of Peierls’ hopping phases consistent with this uniform magnetic flux, the naive particle-hole operation must be composed with a gauge transformation in order to exactly restore all the hopping elements. Given the particular form of the hoppings Eq. (SI) to Eq. (SI) for the two-band model, one may easily confirm that the following form of the particle-hole operation leaves the Hamiltonian invariant:
@@
where $\mathsf{T}$ is the transposition operation on the row vector of electronic annihilation operators (indexed by spin).
To obtain the form of the particle-hole symmetry presented in the main text, we compose the version in Eq. (61) with a conventional spin rotation:
$\mathcal{P}c^{\dagger}_{x,y}\mathcal{P}^{-1}=(-1)^{x+y}c^{\mathsf{T}}_{x,y}(i\sigma_{y}),\quad\mathcal{P}i\mathcal{P}^{-1}=+i.$ (62)
The spin rotation is chosen so that the spin operator
$S_{\alpha}(\bm{r})=\frac{1}{2}c^{\dagger}(\bm{r})\sigma_{\alpha}c(\bm{r})$ (63)
is left invariant. To see this, we note that the unitarity of $\mathcal{P}$ implies that $\mathcal{P}c(\bm{r})\mathcal{P}^{-1}=(-1)^{x+y}(-i\sigma_{y})c^{\dagger\mathsf{T}}(\bm{r})$. This allows us to explicitly compute:
$\mathcal{P}S_{\alpha}(\bm{r})\mathcal{P}^{-1}$ $=\frac{1}{2}\mathcal{P}c^{\dagger}(\bm{r})\mathcal{P}^{-1}\sigma_{\alpha}\mathcal{P}c(\bm{r})\mathcal{P}^{-1}=\frac{1}{2}c^{\mathsf{T}}(\bm{r})\sigma_{y}\sigma_{\alpha}\sigma_{y}c^{\dagger\mathsf{T}}(\bm{r}).$ (64)
When $\alpha\in\{x,z\}$ we acquire a minus sign but then obtain another from fermion anti-commutation when undoing the transposes. On the other hand, when $\alpha=y$ the Pauli matrices commute but the minus sign from fermion anti-commutation cancels with the fact that $\sigma_{y}$ is an anti-symmetric matrix:
$\mathcal{P}S_{y}(\bm{r})\mathcal{P}^{-1}=\frac{1}{2}c_{s}(\bm{r})(\sigma_{y})_{ss^{\prime}}c_{s^{\prime}}^{\dagger}(\bm{r})=-\frac{1}{2}c_{s^{\prime}}^{\dagger}(\bm{r})(\sigma_{y})_{ss^{\prime}}c_{s}(\bm{r})=+S_{y}(\bm{r}).$ (65)
SIII Fate of Symmetries on Cylinder
SIII.1 Particle-hole
Let us elucidate when placing the system on a YC-$L_{y}$ cylinder explicitly breaks particle-hole symmetry. We may use the naive definition of particle-hole Eq. (55) because the fluxes in question are gauge-invariant quantities. Focusing on a particular cylinder ring $r$ penetrated with $\Phi_{r}$ flux, note that
@@
Therefore, $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to a gauge transformation) when, for all rings $r$, we have that
@@
When $L_{y}$ is even, Gauss’s law dictates that the fluxes $\Phi_{r_{1}}$ and $\Phi_{r_{2}}$ through two adjacent rings $r_{1,2}$ must be equal. Note that this allows $T_{x}$ to be a symmetry (up to a gauge transformation). Therefore, possibly after threading $\phi$ external flux through each ring of the cylinder, it is possible to tune all the $\Phi_{r}$ simultaneously (to 0 or $\pi$) so that $\tilde{\mathcal{\mathcal{P}}}$ is respected. Only in those cases is $T_{x}\tilde{\mathcal{\mathcal{P}}}$ also a symmetry (up to gauge).
@@
This always rules out $T_{x}$ as a symmetry. However, so long as $\Phi_{r_{1}}$ is $\pi/2$ or $\pi/2,$ both Eqs. (67) and (68) can be respected, so that $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge). However, odd $L_{y}$ is special in that if we tune $\Phi_{r_{1}}$ to 0 or $\pi,$ then $T_{x}\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge) even though neither $T_{x}$ nor $\tilde{\mathcal{\mathcal{P}}}$ is individually a symmetry.
SIII.2 Magnetic translations
As discussed above, when $L_{y}$ is even, the external magnetic flux through all the plaquettes and loops of the cylinder are consistent with a translation symmetry. In particular, this generally requires composing a bare translation with a gauge transformation. In contrast, the effect of translations on the fluxes through loops cannot be gauged away when $L_{y}$ is odd (see Fig. S1b). Mathematically, this manifests in the fact that $T_{x}$ [see Eq. (49)] is ill-defined, i.e., not single-valued in the coordinate $y$. Thus, any gauge transformation which modifies $T_{x}$ and $T_{y}$ while preserving the property that they anti-commute will also fail to be single-valued.
Appendix D Entanglement spectrum
Figure S3: Entanglement data at $L_{y}=6$ with $\Phi_{\text{ext}}=\pi/2$, taken at the cut to the right of the $x=0$ ring. (a) Weight of the Schmidt values resolved by spin $S_{z}$ and charge $Q$ quantum numbers (see Appendix Sec. D for details) at $U/t=6$. (b) Entanglement eigenvalues vs. circumferential momentum $k_{y}$; colors (with small artificial horizontal separation) indicate spin quantum number (see legend). Left panel: for $Q=0e$, with circled numerals indicating the number of nearly-degenerate $S_{z}=0$ entanglement eigenvalues at each momentum. Right panel: likewise for $Q=1e$, with numerals counting the number of nearly-degenerate $S_{z}=1/2$ states. (c,d) Same for $U/t=14$.
Here we study the real-space entanglement spectrum on infinite-length cylinders at a partition into two semi-infinite halves. The MPS ansatz allows for the efficient computation of the ground state Schmidt decomposition:
$|\Psi\rangle=\sum_{\alpha}\lambda_{\alpha}|\Phi_{\alpha L}\rangle|\Phi_{\alpha R}\rangle,$ (69)
where $|\Phi_{\alpha L/R}\rangle$ are orthonormal bases for the left/right Hilbert spaces. The Schmidt values are non-negative and normalized as $\sum_{\alpha}\lambda_{\alpha}^{2}=1,$ with $\lambda_{\alpha}^{2}$ being the probability of measuring subsystem $R$ in the state $|\Phi_{\alpha R}\rangle$ (which collapses $L$ to $|\Phi_{\alpha L}\rangle$). Tracing out $R$, one obtains a mixed state for $L$ with von Neumann entanglement entropy
$S=-\sum_{\alpha}\lambda_{\alpha}^{2}\log\lambda_{\alpha}^{2}.$ (70)
In Ref. [50], Li and Haldane conjectured that the entanglement spectrum at a bipartition $L\cup R$ of a quantum Hall ground state encodes the low-energy spectrum of the half-system with a physical edge at the location of the bipartition. Given the reduced density matrix $\rho_{L}$, the entanglement spectrum is defined as the eigenvalues of the non-negative-definite Hermitian matrix $H_{L}$ with $\rho_{L}=e^{-H_{L}}$, namely $\epsilon_{\alpha}=-\log\lambda_{\alpha}^{2}$. The Li-Haldane proposal has been further investigated in QH systems [80, 53, 55], demonstrated exactly for free fermions [52], and justified more generally using CFT [54, 56, 57].
The entanglement eigenstates can be labeled by their charge, spin, and circumferential momentum quantum numbers. This enables us to define the Schmidt weight in sector $q=(Q,S_{z},k_{y})$ as the restricted sum $W_{q}=\sum_{\alpha|q}\lambda_{\alpha}^{2}$ over states $\alpha$ with quantum number $q$, so that $\sum_{q}W_{q}=1$. In the infinite geometry, relative charges are unambiguous while the absolute charge is subtle [81]. For concreteness, we use the following convention: we fix the origin $(Q,S_{z})=(0,0)$ so that $Q=0$ has the largest Schmidt weight and the average of $S_{z}$ is zero over the distribution. In each fixed $Q$ sector, we further place the lowest-lying entanglement eigenvalue at $k_{y}=0$ for our plots.
In Fig. S3(a,c) we plot the Schmidt weight resolved by $Q$ and $S_{z}$ for the $L_{y}=6$ cylinder ground states at $U/t=6,14$. We perform the bipartition to the right of ring $x=0$. It is found to be strongly peaked at a single $Q$ for the $U/t=14$ state, but not as a function of $S_{z}$. This indicates weak charge fluctuations (i.e., Mott insulator) and relatively large spin fluctuations. Fig. S3(b), we fix the two most prominent values of $Q$ at $U/t=6$ and plot the ES as a function of $k_{y}$; these spectra carry integer vs. half-integer spin representations, consistent with being related by a unit charge (i.e., an electron). Fig. S3(d) shows the ES for the $U/t=14$ CSL state at $Q=0$, the dominant charge sector. An equivalent ES is obtained at every other cut along the cylinder, indicating the lack of translation SSB.
In Fig. S4 we plot analogous data for the $L_{y}=5$ cylinder with $\Phi_{\text{ext}}=0$, which has the glide-particle-hole symmetry. Within the IQH phase at $U/t=6$, the distribution of Schmidt weights is reflected as a function of $Q$, but nonetheless equivalent, between bond 0 and 1. In contrast, the ES in the $U/t=14$ CSL state differs dramatically at bond 0 vs. 1, with the lowest-lying states carrying half-integer vs. integer spin, indicating SSB of $\mathcal{P}T_{x}$.
Figure S4: Entanglement data at $L_{y}=5$ with $\Phi_{\text{ext}}=0$, analogous to Fig. S3. (a) For the cut to the right of ring $x=0$ (i.e., on bond 0), at $U/t=6$. Entanglement eigenvalues shown in both the $Q=0e$ (upper right panel) and $Q=+1e$ (lower right) sectors. (b) Likewise at bond 1; entanglement spectrum shown for $Q=0e$ and Q=1eQ=-1e sectors. (c,d) Similarly for $U/t=14.$
Appendix E Operators in the hybrid $x\,k_{y}$ representation
In our analysis of the transition, it is useful to study nearest-neighbour hoppings and their correlations. Since the states of interest are translation-invariant, namely around the cylinder circumference, we are motivated to study the ring-averaged correlation functions. For each nearest-neighbour lattice displacement
$\bm{\delta}\in\{\pm\bm{a}_{1},\pm\bm{a}_{2},\pm(\bm{a}_{1}-\bm{a}_{2})\},$ (71)
let us define
$d^{\dagger}_{\bm{\delta}}(x)$ $=\frac{1}{L_{y}}\sum_{y}c^{\dagger}(x,y)c((x,y)+\bm{\delta}).$ (72)
By Eq. (35),
$d^{\dagger}_{\bm{\delta}}(x)$ $=\frac{1}{L_{y}^{2}}\sum_{y}\sum_{k,k^{\prime}}e^{iky}e^{-ik^{\prime}(y+\delta_{y})}c^{\dagger}(x,k)c(x+\delta_{x},k^{\prime})$ (73)
$=\frac{1}{L_{y}^{2}}\sum_{y}\sum_{k,k^{\prime}}e^{i(k-k^{\prime})y}e^{-ik^{\prime}\delta_{y}}c^{\dagger}(x,k)c(x+\delta_{x},k^{\prime})$ (74)
$=\frac{1}{L_{y}}\sum_{k}e^{-ik\delta_{y}}c^{\dagger}(x,k)c(x+\delta_{x},k),$ (75)
so that
$d^{\dagger}_{\bm{a}_{1}}(x)$ $=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x+1,k)$ (76)
$d^{\dagger}_{-\bm{a}_{1}}(x)$ $=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x-1,k)$ (77)
$d^{\dagger}_{\bm{a}_{2}}(x)$ $=\frac{1}{L_{y}}\sum_{k}e^{-ik}c^{\dagger}(x,k)c(x,k)$ (78)
$d^{\dagger}_{-\bm{a}_{2}}(x)$ $=\frac{1}{L_{y}}\sum_{k}e^{+ik}c^{\dagger}(x,k)c(x,k),$ (79)
and similarly for $\bm{\delta}=\pm(\bm{a}_{1}-\bm{a}_{2})$.
SI Staggered Heisenberg Order Parameter
The following documents the numerical computation of the staggered Heisenberg order parameter in the hybrid $xk_{y}$ fermion representation. The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can be expressed in terms of fermionic operators by first noting that
$S_{i}^{+}=c_{i\uparrow}^{\dagger}c_{i\downarrow},\quad S_{i}^{-}=c_{i\downarrow}^{\dagger}c_{i\uparrow},\quad S_{i}^{z}=\frac{1}{2}(c_{i\uparrow}^{\dagger}c_{i\uparrow}-c_{i\downarrow}^{\dagger}c_{i\downarrow}).$ (80)